# Why E.T. Can’t Phone Home: A Global View on IP-based Geoblocking at VoWiFi

### v1

@@ Abstract.
This work evaluates the current deployment status of VoWiFi among worldwide operators and analyzes existing geoblocking measures on the IP layer. We show that a substantial share (IPv4: 14.6%, IPv6: 65.2%) of operators implement geoblocking at the DNS- or VoWiFi protocol level, and highlight severe drawbacks in terms of emergency calling service availability.
@@
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–07, 2024; Tokyo, Japanbooktitle: MobiSys ’24: The 22nd Annual International Conference on Mobile Systems, Applications and Services (MobiSys ’24), June 03–07, 2023, Tokyo, Japanprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Networks Mobile networksccs: Security and privacy Mobile and wireless securityccs: Networks Network management
@@ 1. Introduction
Mobile network services are a crucial lifeline in today’s society, given that in 2023 over 5.4 billion people rely on cellular networks for connectivity and communication (GSMA2023MobileEconomy, 42). With 4G currently being the most used wireless standard and 5G rapidly gaining penetration, numerous operators are actively decommissioning older legacy networks (2G and 3G), marking the completion of the shift from circuit-switched to a comprehensive packet-switched network paradigm.
In the packet-switched domain, operators use VoIP (Voice over IP) based technology to terminate voice calls and messages. Additionally to the VoLTE (Voice over LTE) standard, VoWiFi (Voice over Wi-Fi, also known as Wi-Fi calling) was introduced. While VoLTE uses the traditional radio infrastructure that is provided by the operator as its access medium, VoWiFi is a complementary solution that allows the use of third-party wireless networks as an alternative uplink to the operator. Consequently, customers can leverage existing Wi-Fi access points and continue utilizing their mobile phones for voice calls in areas with poor or no cellular reception.
Figure 1. An Indian operator states in the FAQs (Airtel2023Blocking, 2) that VoWiFi cannot be used during international roaming.
@@ 2. Background
3GPP Access, Voice over LTE (VoLTE). As shown in the upper path of Figure 2, the User Equipment (UE) attaches to a base station (Evolved Node B (eNB) in LTE) via the Radio Acces Network (RAN). The Serving Gateway (S-GW) and the Packet Data Network Gateway (P-GW) within the Evolved Packet Core (EPC) system are responsible for assigning IP addresses to Access Point Names (APNs) and for routing and forwarding the traffic to external Packet Data Networks (PDNs). Finally, the Proxy Call Session Control Function (P-CSCF) acts as a Session Initiation Protocol (SIP) proxy and is the ingress point to the IP Multimedia Subsystem (IMS). All data traffic between the UE and the P-CSCF is encapsulated in an IPsec tunnel. The UE can then directly send SIP messages, e.g., after successful establishment of the IPSec tunnel it can send a SIP REGISTER request to the IMS core network. While SIP is used for signaling, the actual audio stream of a call is transferred via the Real-Time Transport Protocol (RTP) protocol.
@@
Internet Protocol Security (IPsec). As described above, VoLTE and VoWiFi heavily rely on IPSec (rfc6071, 16) for authentication and traffic encapsulation. To set up a Security Association (SA) it uses the Internet Key Exchange (IKE) protocol. More specifically it uses IKEv2 (rfc4306, 24, 25) with EAP-AKA (rfc4187, 4) (Authentication and Key Agreement) for key derivation and thereby leverages the SIM card’s secret keys to obtain a new session key. The negotiation can be divided into two phases: IKE_SA_INIT that negotiates the ciphering suite and other security parameters, and IKE_AUTH where the SIM card authenticates by solving a random challenge using its secret keys.
@@ 3.1. Mapping DNS Records
The Fully Qualified Domain Name (FQDN) of an ePDG is specified in 3GPP TS 23.003 (EtsiNumberingAdressingIdentification, 12) and can be built from an operator’s Mobile Country Code (MCC) and Mobile Network Code (MNC):
@@
If we want to get an exhaustive list of ePDGs from all operators around the globe, requesting the IP addresses via normal (recursive) DNS requests (i.e., offloading them to popular DNS servers like Cloudflare or Google) from one central location does not work, or would at least yield imprecise or non-deterministic results (e.g., due to caching, Anycast routing, etc.). To make recursive DNS servers query for geographically close IP addresses, the eDNS Client Subnet (ECS) (rfc7871, 7) mechanism allows propagating the client’s IP address range (usually a /24 subnet) to the authoritative DNS server. However, some DNS servers (e.g., Cloudflare (CloudflareFAQECS, 6)) do not enable ECS due to privacy concerns. Also, we found several operators’ authoritative DNS servers do not support ECS and therefore solely use the request’s IP address as baseline to build their responses. Addressing this requires a more complex approach, that uses different locations worldwide to issue DNS requests in an authoritative manner.
@@
NordVPN 60 $\times$
Private Internet Access 84 $\times$
ProtonVPN 68 $\times$
Surfshark 100 $\times$
${}^{\mathrm{a}}$ As advertised by the VPN/cloud service.
@@ 3.2. Discovering DNS Records
Region (via MCC) ${}^{\mathrm{a}}$ Countries ${}^{\mathrm{b}}$ Domains IPs Countries ${}^{\mathrm{b}}$ Domains IPs
@@ 9 Worldwide ${}^{\mathrm{c}}$ 1 1 1 0 0 0
${}^{\mathrm{a}}$ digits 1 and 8 are not specified.   ${}^{\mathrm{b}}$ according to the MCC.   ${}^{\mathrm{c}}$ Satellite, Air, Maritime, Antarctica.
Table 2. Encountered ePDGs via DNS discovery, grouped by MCC region.
@@ 4. Results and Evaluation
When citing a particular operator, we employ the notation CarrierName${}_{[MCCMNC]}$.
@@
The vast majority of providers use three digits for the MNC in their ePDG domain, while Vodacom${}_{[64004]}$ (Tanzania) is the only one that reserves an additional two-digit domain (i.e., epdg.epc.mnc04.mcc640.pub.3gppnetwork.org), serving the same IPs as its three digit counterpart.
Geographic Location of ePDGs. To determine the country of origin for an MNO (by its MCC), we rely on the most recent version of Android’s MCC table (AOSPMCC, 1), which is based on T-SP-E.212A (ITUT, 23) standardized by the International Telecommunication Union (ITU). To geolocate IP addresses we use the free MaxMind GeoLite2 database [^5] https://dev.maxmind.com/geoip/geolite2-free-geolocation-data. The MCC country of an ePDG and the location of an ePDG according to its IP addresses do not necessarily need to be identical. However, the majority of the operators use ePDGs that are hosted within their own network range and country. In most cases where the ePDG is not located within the MCC country, it is hosted within close proximity (i.e., in a neighbouring country). We noticed that this practice is especially popular with relatively small countries (e.g., Tele2${}_{[24603,24702]}$ in Latvia and Lithuania via Sweden, Orange${}_{[27099]}$ in Luxembourg via Belgium or Claro${}_{[74810]}$ in Uruguay via Argentina) and in Caribbean island countries (e.g., Flow${}_{[365840,364390]}$ in Anguilla and the Bahamas via Jamaica). An outlier with no geographical proximity of MCC and ePDG country is Tata Communications${}_{[23427]}$ which is based in the United Kingdom but uses an ePDG located in the United States. MTX Connect${}_{[90139]}$, the only operator we found within the “Worldwide” MCC range, is pointing to an ePDG IP hosted in Luxembourg. For IPv6 we do not see any divergences between MCC and ePDG country.
Non-Routable IP Addresses. While the majority of returned IP addresses lie within public address ranges, some results are not publicly routable. For example, German Voiceworks${}_{[26220]}$ and Italian Wind Tre${}_{[22288,22299]}$ return loopback IP addresses (127.0.0.1 and 127.0.0.9). Obviously, these addresses will not work in practice and were presumably just deployed as a placeholder or for internal testing purposes.
For IPv6, we see several providers referring to their IPv4 siblings. More specifically, Three Mobile${}_{[23594]}$ (United Kingdom) and Méditélécom${}_{[60400]}$ (Morocco) use the NAT64 IPv6 transition mechanism (rfc6146, 36) via the 64:ff9b::/96 prefix and Maxis${}_{[50212]}$ (Malaysia) refers to its IPv4 sibling via an IPv4-Compatible IPv6 address (deprecated in RFC4291 (rfc4291, 9)).
@@ 4.1. Analyzing Differences in DNS Responses
Table 3. To achieve global coverage and to improve diversity of our client-side vantage points, we evenly distributed the IKE probing measurements to numerous VPN- and cloud services. Surprisingly, Cloudflare Warp provided far more countries than advertised in the service description (120 advertised vs. 208 actual countries).
@@
    Returning Multiple IPs for Redundancy.  This group contains all domains where the DNS server directly returns all IP addresses that are associated with the queried hostname (without differentiating by the client’s location). The vast majority of all domains (at least 78% [^6] According to our simple heuristic algorithm. of all IPv4 domains) show this behavior. Responding with a greater set of IP addresses makes sense from an availability perspective: a client that wants to connect to a service can always switch to another IP address if something goes wrong when connecting to the first endpoint. Additionally, due to round-robin DNS (rfc1794, 5), this will also enforce automatic load-balancing across all associated IP addresses.
    While the median number of IP addresses that are returned is only two, some operators return a greater number of IPs. For instance, the maximum number that we discovered was Telekom${}_{[26201]}$ (Germany) where all 12 IPs that exist for their ePDG domain are instantly returned by the responsible DNS server.
@@
    T-Mobile${}_{[310240,310260]}$ (United States) stands out as the primary example within this category, as they provide a total of 39 IP addresses for each of their domains. All ePDG servers were in deployment concurrently (i.e., discovered during various scans in the same time period) but their DNS server only answers with a single IP per request, as described above. Supposedly, the returned IP is selected randomly, as we do not see any location bias.
@@
    Using DNS for Geolocational Grouping.  Again, only a single entry of a bigger address pool is returned. Additionally, the returned address is determined by the source IP address. Analyzing the IPs returned for specific countries, we see that the operator uses predetermined IP addresses to serve customers in different locations. For example, Reliance Jio${}_{[405874]}$ [^7] Jio also uses 21 other MNCs that were shortened for the sake of visibility. (India) clearly separates their domestic and foreign users. All queries issued within India received IP addresses that never occurred within any other country. For their external customers, we could not see any further differentiation (i.e., customers in Europe usually get the same responses as customers in America or Africa). Additionally, we noticed that their DNS server does not support the eDNS Client Subnet (ECS) (rfc7871, 7) mechanism. Therefore, the only way to discover their local IP addresses is to issue DNS queries from a location within India.
@@
    Using DNS for Geoblocking.  Finally, operators could only answer local DNS queries and simply block or drop any external requests, to prohibit their customers from accessing the IMS via Wi-Fi when being abroad. In our results, we found that Vodafone${}_{[26202]}$ (Germany) is only serving its ePDG IP addresses for DNS requests from appropriate client IPs. This was occasionally noticed by actual customers, as we also found anecdotal evidence (VodafoneVoWiFiIssue1, 3, 51, 53, 43, 21, 50) within several posts on blogs and online forums. In contrast to Jio${}_{[405874]}$ (see above), Vodafone${}_{[26202]}$ actually respects ECS queries, making it possible to easily demonstrate the filtering [^8] Example queries are shown in the Appendix A.1. While most of the queries that were able to discover Vodafone${}_{[26202]}$’s ePDG were issued within Germany, the DNS server occasionally answered requests from external countries (e.g., several close countries like Austria, and Ireland, but also more distant countries like Kazakhstan and Japan). We tried to investigate what caused these false positives and checked the IP information of several results according to the MaxMind database. We found, that many misclassified IPs had Germany set as the registered_country (an additional meta-information in the database), although delivering an external country as the primary hosting country.
IPv6. Again, the vast majority (over 90 %) of all domains that return AAAA records can be found within the first group G1. Additionally, we found members of G2 and G3, namely Israeli Cellcom${}_{[42502]}$, replying with a single IP chosen without any visible location bias ( G2) during a transition period switching their ePDG server, and Canadian Bell${}_{[302610]}$, that serve their customers with dedicated IPs that are (vaguely) grouped by geographic region. In contrast to IPv4, we did not see any operator that used DNS for geoblocking purposes.
@@ 4.2. Probing VoWiFi Servers
Data Sources and Covered Countries. Both MaxMind (maxmind, 37) and Android’s MCC table (AOSPMCC, 1) link entities (i.e., IP addresses, operators) to a geographical region (i.e., an ISO 3166-1 (ISO3166-1, 22) country). ISO 3166-1 currently comprises 249 countries, and each can be linked to its corresponding continent.

### v2

@@ Abstract.
This work evaluates the current deployment status of VoWiFi among worldwide operators and analyzes existing geoblocking measures on the IP layer by measuring connectivity from over 200 countries. We show that a substantial share (IPv4: 14.6%, IPv6: 65.2%) of operators implement geoblocking at the DNS- or VoWiFi protocol level, and highlight severe drawbacks in terms of emergency calling service availability.
@@
copyright: rightsretainedjournalyear: 2024conference: The 22nd Annual International Conference on Mobile Systems, Applications and Services; June 3–7, 2024; Minato-ku, Tokyo, Japanbooktitle: The 22nd Annual International Conference on Mobile Systems, Applications and Services (MOBISYS ’24), June 3–7, 2024, Minato-ku, Tokyo, Japandoi: 10.1145/3643832.3661883isbn: 979-8-4007-0581-6/24/06ccs: Networks Mobile networksccs: Security and privacy Mobile and wireless securityccs: Networks Network management
@@ 1. Introduction
Mobile network services are a crucial lifeline in today’s society, given that in 2023 over 5.4 billion people relied on cellular networks for connectivity and communication (GSMA2023MobileEconomy). With 4G currently being the most used wireless standard and 5G rapidly gaining penetration, numerous operators are actively decommissioning older legacy networks (2G and 3G), marking the completion of the shift from circuit-switched to a comprehensive packet-switched network paradigm.
In the packet-switched domain, operators use VoIP (Voice over IP) based technology to terminate voice calls and messages. Additionally to the VoLTE (Voice over LTE) standard, VoWiFi (Voice over Wi-Fi, also known as Wi-Fi calling) was introduced. While VoLTE uses the traditional radio infrastructure that is provided by the operator as its access medium, VoWiFi is a complementary solution that allows the use of third-party wireless networks as an alternative uplink to the operator. Consequently, customers can leverage existing Wi-Fi access points (APs) and continue utilizing their mobile phones for voice calls in areas with poor or no cellular reception.
Figure 1. An Indian operator states in the FAQs (Airtel2023Blocking) that VoWiFi cannot be used during international roaming.
@@ 2. Background
3GPP Access, Voice over LTE (VoLTE). As shown in the upper path of Figure 2, the User Equipment (UE) attaches to a base station (Evolved Node B (eNB) in LTE) via the Radio Acces Network (RAN). The Serving Gateway (S-GW) and the Packet Data Network Gateway (P-GW) within the Evolved Packet Core (EPC) system are responsible for assigning IP addresses to Access Point Names (APNs) and for routing and forwarding the traffic to external Packet Data Networks (PDNs). Finally, the Proxy Call Session Control Function (P-CSCF) acts as a Session Initiation Protocol (SIP) proxy and is the ingress point to the IP Multimedia Subsystem (IMS). All data traffic between the UE and the P-CSCF is encapsulated in an IPsec tunnel. The UE can then directly send SIP messages, e.g., after successful establishment of the IPSec tunnel it can send a SIP REGISTER request to the IMS core network. While SIP is used for signaling, the actual audio stream of a call is transferred via the Real-Time Transport Protocol (RTP).
@@
Internet Protocol Security (IPsec). As described above, VoLTE and VoWiFi heavily rely on IPSec (rfc6071) for authentication and traffic encapsulation. To set up a Security Association (SA) it uses the Internet Key Exchange (IKE) protocol. More specifically it uses IKEv2 (rfc4306; rfc7296) with EAP-AKA (rfc4187) (Authentication and Key Agreement) for key derivation and thereby leverages the SIM card’s secret keys to obtain a new session key. The negotiation can be divided into two phases: IKE_SA_INIT that negotiates the ciphering suite and other security parameters, and IKE_AUTH where the SIM card authenticates by solving a random challenge using its secret keys.
@@ 3.1. Mapping DNS Records
The Fully Qualified Domain Name (FQDN) of an ePDG is specified in 3GPP TS 23.003 (EtsiNumberingAdressingIdentification) and can be built from an operator’s Mobile Country Code (MCC) and Mobile Network Code (MNC):
@@
If we want to get an exhaustive list of ePDGs from all operators around the globe, requesting the IP addresses via normal (recursive) DNS requests (i.e., offloading them to popular DNS servers like Cloudflare or Google) from one central location does not work, or would at least yield imprecise or non-deterministic results (e.g., due to caching, Anycast routing, etc.). To make recursive DNS servers query for geographically close IP addresses, the eDNS Client Subnet (ECS) (rfc7871) mechanism allows propagating the client’s IP address range (usually a /24 subnet) to the authoritative DNS server. However, some DNS servers (e.g., Cloudflare (CloudflareFAQECS)) do not enable ECS due to privacy concerns. Also, we found several operators’ authoritative DNS servers do not support ECS and therefore solely use the request’s IP address as baseline to build their responses. Addressing this requires a more complex approach, that uses different locations worldwide to issue DNS requests in an authoritative manner.
@@
NordVPN 60 $\times$×
Private Internet Access 84 $\times$×
ProtonVPN 68 $\times$×
Surfshark 100 $\times$×
a As advertised by the VPN/cloud service.
@@ 3.2. Discovering DNS Records
Region (via MCC) a Countries b Domains IPs Countries b Domains IPs
@@ 9 Worldwide c 1 1 1 0 0 0
a digits 1 and 8 are not specified.   b according to the MCC.   c Satellite, Air, Maritime, Antarctica.
Table 2. Encountered ePDGs via DNS discovery, grouped by MCC region. For Test networks we found no public DNS entries.
@@ 4. Results and Evaluation
When citing a particular operator, we employ the notation CarrierName[MCCMNC].
@@
The vast majority of providers use three digits for the MNC in their ePDG domain, while Vodacom[64004] (Tanzania) is the only one that reserves an additional two-digit domain (i.e., epdg.epc.mnc04.mcc640.pub.3gppnetwork.org), serving the same IPs as its three digit counterpart.
Geographic Location of ePDGs. To determine the country of origin for an MNO (by its MCC), we rely on the most recent version of Android’s MCC table (AOSPMCC), which is based on T-SP-E.212A (ITUT) standardized by the International Telecommunication Union (ITU). To geolocate IP addresses we use the free MaxMind GeoLite2 database [^5] https://dev.maxmind.com/geoip/geolite2-free-geolocation-data. The MCC country of an ePDG and the location of an ePDG according to its IP addresses do not necessarily need to be identical. However, the majority of the operators use ePDGs that are hosted within their own network range and country. In most cases where the ePDG is not located within the MCC country, it is hosted within close proximity (i.e., in a neighbouring country). We noticed that this practice is especially popular with relatively small countries (e.g., Tele2[24603,24702] in Latvia and Lithuania via Sweden, Orange[27099] in Luxembourg via Belgium or Claro[74810] in Uruguay via Argentina) and in Caribbean island countries (e.g., Flow[365840,364390] in Anguilla and the Bahamas via Jamaica). An outlier with no geographical proximity of MCC and ePDG country is Tata Communications[23427] which is based in the United Kingdom but uses an ePDG located in the United States. MTX Connect[90139], the only operator we found within the “Worldwide” MCC range, is pointing to an ePDG IP hosted in Luxembourg. For IPv6 we do not see any divergences between MCC and ePDG country.
Non-Routable IP Addresses. While the majority of returned IP addresses lie within public address ranges, some results are not publicly routable. For example, German Voiceworks[26220] and Italian Wind Tre[22288,22299] return loopback IP addresses (127.0.0.1 and 127.0.0.9). Obviously, these addresses will not work in practice and were presumably just deployed as a placeholder or for internal testing purposes.
For IPv6, we see several providers referring to their IPv4 siblings. More specifically, Three Mobile[23594] (United Kingdom) and Méditélécom[60400] (Morocco) use the NAT64 IPv6 transition mechanism (rfc6146) via the 64:ff9b::/96 prefix and Maxis[50212] (Malaysia) refers to its IPv4 sibling via an IPv4-Compatible IPv6 address (deprecated in RFC4291 (rfc4291)).
@@ 4.1. Analyzing Differences in DNS Responses
Table 3. To achieve global coverage and to improve diversity of our client-side vantage points, we evenly distributed the IKE probing measurements to numerous VPN- and cloud services. Surprisingly, Cloudflare Warp provided far more countries than advertised in the service description (120 advertised vs. 208 actual countries). While they only list locations of their data centers the service presumably also uses smaller edge locations as exit points.
@@
    Returning Multiple IPs for Redundancy.  This group contains all domains where the DNS server directly returns all IP addresses that are associated with the queried hostname (without differentiating by the client’s location). The vast majority of all domains (at least 78% [^6] According to our simple heuristic algorithm. of all IPv4 domains) show this behavior. Responding with a greater set of IP addresses makes sense from an availability perspective: a client that wants to connect to a service can always switch to another IP address if something goes wrong when connecting to the first endpoint. Additionally, due to round-robin DNS (rfc1794), this will also enforce automatic load-balancing across all associated IP addresses.
    While the median number of IP addresses that are returned is only two, some operators return a greater number of IPs. For instance, the maximum number that we discovered was Telekom[26201] (Germany) where all 12 IPs that exist for their ePDG domain are instantly returned by the responsible DNS server.
@@
    T-Mobile[310240,310260] (United States) stands out as the primary example within this category, as they provide a total of 39 IP addresses for each of their domains. All ePDG servers were in deployment concurrently (i.e., discovered during various scans in the same time period) but their DNS server only answers with a single IP per request, as described above. Supposedly, the returned IP is selected randomly, as we do not see any location bias.
@@
    Using DNS for Geolocational Grouping.  Again, only a single entry of a bigger address pool is returned. Additionally, the returned address is determined by the source IP address. Analyzing the IPs returned for specific countries, we see that the operator uses predetermined IP addresses to serve customers in different locations. For example, Reliance Jio[405874] [^7] Jio also uses 21 other MNCs that were shortened for the sake of visibility. (India) clearly separates their domestic and foreign users. All queries issued within India received IP addresses that never occurred within any other country. For their external customers, we could not see any further differentiation (i.e., customers in Europe usually get the same responses as customers in America or Africa). Additionally, we noticed that their DNS server does not support the eDNS Client Subnet (ECS) (rfc7871) mechanism. Therefore, the only way to discover their local IP addresses is to issue DNS queries from a location within India.
@@
    Using DNS for Geoblocking.  Finally, operators could only answer local DNS queries and simply block or drop any external requests, to prohibit their customers from accessing the IMS via Wi-Fi when being abroad. In our results, we found that Vodafone[26202] (Germany) is only serving its ePDG IP addresses for DNS requests from appropriate client IPs. This was occasionally noticed by actual customers, as we also found anecdotal evidence (VodafoneVoWiFiIssue1; VodafoneVoWiFiIssue2; VodafoneVoWiFiIssue3; VodafoneVoWiFiIssue4; VodafoneVoWiFiIssue5; VodafoneVoWiFiIssue6) within several posts on blogs and online forums. In contrast to Jio[405874] (see above), Vodafone[26202] actually respects ECS queries, making it possible to easily demonstrate the filtering [^8] Example queries are shown in the Appendix A.1. While most of the queries that were able to discover Vodafone[26202]’s ePDG were issued within Germany, the DNS server occasionally answered requests from external countries (e.g., several close countries like Austria, and Ireland, but also more distant countries like Kazakhstan and Japan). We tried to investigate what caused these false positives and checked the IP information of several results according to the MaxMind database. We found, that many misclassified IPs had Germany set as the registered_country (an additional meta-information in the database), although delivering an external country as the primary hosting country.
IPv6. Again, the vast majority (over 90 %) of all domains that return AAAA records can be found within the first group G1. Additionally, we found members of G2 and G3, namely Israeli Cellcom[42502], replying with a single IP chosen without any visible location bias ( G2) during a transition period switching their ePDG server, and Canadian Bell[302610], that serve their customers with dedicated IPs that are (vaguely) grouped by geographic region. In contrast to IPv4, we did not see any operator that used DNS for geoblocking purposes.
@@ 4.2. Probing VoWiFi Servers
Data Sources and Covered Countries. Both MaxMind (maxmind) and Android’s MCC table (AOSPMCC) link entities (i.e., IP addresses, operators) to a geographical region (i.e., an ISO 3166-1 (ISO3166-1) country). ISO 3166-1 currently comprises 249 countries, and each can be linked to its corresponding continent.
@@
The number of countries we used for scanning greatly exceeds the number of countries where operators actually support and use VoWiFi in practice. To maximize the scope of our study and gain a defacto worldwide view, we scanned from all 219 available countries. More specifically, this was done i) to maximize the discovered DNS entries in case of DNS-based blocking, and ii) to also detect blocking of countries that do not have VoWiFi yet, but are blocked by foreign operators (e.g., for political reasons).